(3.1) Find an expression for 1

2

||~u + ~v||2 +

1

2

||~u − ~v||2

in terms of ||~u||2 + ||~v||2

.

(3.2) Find an expression for ||~u + ~v|| 2 − ||~u − ~v||2

in terms of ~u · ~v

(3.3) Use the result of (3.2) to deduce an expression for ||~u + ~v||2 whenever ~u and ~v are orthogonal

to each other.


1
Expert's answer
2021-06-23T17:37:22-0400

(3.1)


"(\vec u+\vec v)^2=(\vec u+\vec v)\cdot(\vec u+\vec v)"

"=(\vec u, \vec u)+2(\vec u, \vec v)+(\vec v, \vec v)"

"=||\vec u||^2+2(\vec u, \vec v)+||\vec v||^2"


"(\vec u-\vec v)^2=(\vec u-\vec v)\cdot(\vec u-\vec v)"

"=(\vec u, \vec u)-2(\vec u, \vec v)+(\vec v, \vec v)"

"=||\vec u||^2-2(\vec u, \vec v)+||\vec v||^2"




"\dfrac{1}{2}||\vec u+\vec v||^2+\dfrac{1}{2}||\vec u-\vec v||^2"

"=\dfrac{1}{2}(||\vec u||^2+2(\vec u, \vec v)+||\vec v||^2)+\dfrac{1}{2}(||\vec u||^2-2(\vec u, \vec v)+||\vec v||^2)"

"=||\vec u||^2+||\vec v||^2"

(3.2)


"||\vec u+\vec v||^2-||\vec u-\vec v||^2"

"=||\vec u||^2+2(\vec u, \vec v)+||\vec v||^2-(||\vec u||^2-2(\vec u, \vec v)+||\vec v||^2)"

"=4(\vec u, \vec v)"

(3.3)

If "\vec u \perp \vec v," then "(\vec u, \vec v)=0"


"||\vec u+\vec v||^2=||\vec u||^2+||\vec v||^2=||\vec u-\vec v||^2"